# WildGraph: Realistic Graph-based Trajectory Generation for Wildlife

### v1

@@ 5.5. Results
Table 1. Test Results for the geese and stork datasets using WildGraph and other benchmarks using different measurement
Dataset Method Hausdorff DTW FDE r-Coefficient $chi^{2}$
val cov val cov val cov
Geese VAE 7.45 0.57 57.67 0.68 3.81 0.80 0.47 1588
GAN 25.25 0.30 196.8 0.27 11.87 0.44 -0.47 11894
WildGEN 7.12 0.72 48.60 0.79 3.2 0.86 0.79 772
WildGraph 9.38 0.93 67.87 0.88 2.61 0.78 0.79 725
WildGraph-Transformer 22.01 0.90 225.73 0.85 14.26 0.82 0.53 2232
Stork VAE 7.74 0.64 41.1 0.58 3.86 0.53 0.56 676
GAN 14.54 0.19 58.97 0.17 3.96 0.34 0.61 1501
WildGEN 5.55 0.61 28.41 0.58 3.12 0.61 0.76 611
WildGraph 7.29 0.80 40.16 0.77 3.74 0.69 0.8 357
WildGraph-Transformer 16.07 0.64 92.92 0.62 9.45 0.51 0.62 631
Table 1 lists the results obtained for WildGraph and the other benchmarks on the metrics described above. WildGraph achieves superior overall performance on most generalization measurements, including coverage results, r-Coeff, and $chi^{2}$. On the distance metrics, WildGraph reports comparable results to VAE and WildGEN on both datasets.
5.5.1. Split Threshold
As discussed in section 4.1, here, we analyze the affect of the split threshold on generalization. Figure 6 depicts how coverage values, and r-Coeff is affected for different values of the split threshold.
Figure 6. Generalization as a function of split threshold
Based on the results of Figure 6, a split threshold (r) value of 0.25 is used in all of our experiments, as it provides the highest r-Coeff.
@@ 5.6. Ablation Study
In the ablation study, we want to measure the effects of the HNG and the node2vec embeddings on the measured metrics. This helps explain how each component contributes to the overall structure.
Table 2. Ablation Study on the Geese Dataset
Method Haus. DTW FDE r-Coef
val cov val cov val cov
Uniform H3 / Coarse 9.65 0.85 74.1 0.80 3.90 0.88 0.63
Uniform H3 / Granular 10.58 0.83 84.49 0.87 3.47 0.67 0.75
BOW (no node2vec) 37.5 0.08 224.16 0.08 0.74 0.17 0.73
No PE 9.06 0.88 69.46 0.83 3.6 0.80 0.72
WildGraph 9.38 0.80 67.87 0.88 2.61 0.78 0.79
@@
As evident from Table 2, node embeddings emerge as the most impactful component within the WildGraph framework based on the observed results. Using BOW instead provided poor results possibly due to both its high dimensionality and the absence of node relationships in the embeddings.
5.7. Qualitative Analysis
Figure 7 depicts the performance of WildGraph in comparison to other benchmarks. A preliminary visual examination reveals that WildGraph trajectories exhibit more realistic paths. Unlike WildGEN and VAEs, and similar to the real samples, trajectories do not cluster in the middle of the sea. It is also observed that GAN and Transformer-based approaches provided poor results, which is expected given the small data nature of the problem.
@@ 6. Conclusions
In this paper, we study the problem of trajectory generation over a long horizon, with a focus on applications in wildlife in a small data setting. Existing methods fail to consider the underlying network that describes that movement. In order to solve this problem, we introduce a hierarchical prototype network graph that captures the global movement characteristics and recursively localizes regions. We propose a novel Variational Recurrent Network to capture the transitions and feed them into a latent dictionary. The effective use of the H3 library, and node2vec embeddings has contributed to the success of the WildGraph framework, as measured in the experimental results and ablation study.
This study adds to the body of literature where graph based solutions can effectively address unique problems. Ours is the first one to show how one can generate trajectories for wildlife movement. In our case, the small data nature of wildlife datasets combined with wide horizon predictions presented significant challenges. Our results demonstrate the effectiveness of our proposed approaches that utilize specially designed encoder-decoder networks. In the near future, we will explore how our generated trajectories can be used for data augmentation and training larger scale models in small data settings.

### v2

@@ 5.6. Results
Table 1 lists the results obtained for WildGraph and the other benchmarks on the abovementioned metrics. WildGraph achieves superior overall performance on most coverage and likeness measurements. Regarding the distance metrics, WildGraph reports results comparable to those of VAE and WildGEN on both datasets, with some exceptions where WildGEN outperforms WildGraph. However, given that WildGraph provides better coverage results, which implies that the generated trajectories from WildGraph generalize better on the held-out set. On the other hand, WildGEN fails to generalize on a larger subset of the held-out set compared to WildGraph. A key limitation of WildGEN is that it applies to a continuous spatial domain, leading to unrealistic paths, such as birds flying across the ocean. By discretizing the spatial domain, WildGraph ensures generated paths remain within realistic regions. This makes WildGraph more useful for generating trajectories similar to real, unknown paths.
Similarly, while WildGEN provides a higher SatCLIP result, it also fails to generalize as effectively as WildGraph. WildGraph outperforms all other methods on this measure. However, this finding also suggests that while paths traversed by birds can be captured well to some extent using SatCLIP, it may fail to address all the nuances of the movement. To understand this further, we measure the average pairwise similarity between real data as 0.4845 and 0.4805 for the geese and stork datasets, respectively. In the experiments, however, the results considered only the closest matches from the training set against each trajectory in the held-out set, as discussed above, which may explain the higher scores.
In summary, the results demonstrate WildGraph’s exceptional capacity to learn the properties of very small datasets but still generate new trajectories in a meaningful way that exceeds all other methods in terms of generalization while providing comparable distance, likeness, and geographic similarity measures.
@@ 5.7. Ablation Study
Table 3. Test Results for the geese and stork datasets that demonstrate the contribution of key components and constructs
Dataset Method Hausdorff DTW FDE SatCLIP r-Coefficient $chi^{2}$
value cov value cov value cov value cov
Geese Uniform H3 / Coarse 10.02 0.93 73.68 0.92 4.03 0.87 0.51 0.85 0.62 815
Uniform H3 / Granular 10.3 0.85 83.18 0.82 3.21 0.75 0.56 0.8 0.68 889
BoW (no node2vec) 15.99 0.1 108.11 0.08 1.67 0.17 0.54 0.1 0.43 8364
No PE 8.07 0.92 59.63 0.9 3.02 0.82 0.53 0.93 0.71 762
WildGraph 9.63 0.93 71.08 0.92 2.93 0.81 0.57 0.89 0.80 628
Stork Uniform H3 / Coarse 6 0.84 32.57 0.77 3.07 0.75 0.58 0.81 0.7 339
Uniform H3 / Granular 7.27 0.57 41.84 0.6 4.41 0.56 0.6 0.75 0.78 348
BoW (no node2vec) 12.36 0.13 52.39 0.08 2.02 0.08 0.49 0.09 0.13 3932
No PE 6.57 0.79 38.18 0.74 4.14 0.67 0.59 0.77 0.73 421
WildGraph 7.11 0.84 38.11 0.8 3.67 0.74 0.59 0.85 0.74 369
We perform two ablation studies:
  1. (1)
    Measure how each component of WildGraph contributes to the results achieved.
  2. (2)
    Investigate how different values of the split threshold affect the results.
As evident from Table 3, node embeddings are the most impactful component of the WildGraph framework based on the observed results. Using BoW, however, provided poor results, possibly due to its high dimensionality and the lack of node relationships in the embeddings. Additionally, PE significantly enhances the coverage of the methods. Dynamically breaking the regions with the HNG significantly improves likeness, while using coarse or granular regions shows limited effectiveness. These results highlight WildGraph’s robustness and efficiency across different evaluation metrics when combining all its components.
For additional insight into our solution, we also evaluate how different values of split threshold impact the results. This is summarized in Table 4. Subtle variations in the results as a function of the split threshold is observed. The choice of threshold can be motivated by the dataset at hand. While WildGraph (using the specified split-threshold) is not the best in every metric, it’s close to the best in both distance and coverage. No method outperforms all others in every metric. On the whole, we believe a threshold of $1.0$ is the best overall and provides the best compromise on these metrics for both datasets. However, for trajectories where the length is shorter, a smaller threshold may provide a better outcome.
5.8. Qualitative Analysis
Figure 7 depicts the performance of WildGraph in comparison to other benchmarks. A preliminary visual examination reveals that WildGraph trajectories exhibit more realistic paths. Unlike WildGEN and VAEs, and similar to the real samples, trajectories do not cluster in the middle of the sea. It is also observed that GAN and Transformer-based approaches provided poor results, which is expected given the small data nature of the problem. More results including Levy Flight are omitted for brevity and may be examined in (Al-Lawati et al., 2023).
Table 4. Test Results for the geese and stork datasets using WildGraph with different split threshold configurations
Dataset Split Threshold Hausdorff DTW FDE SatCLIP r-Coefficient $chi^{2}$
value cov value cov value cov value cov
Geese 0.25 9.55 0.93 71.28 0.9 2.76 0.79 0.57 0.87 0.8 614
0.5 9.62 0.92 70.06 0.91 2.96 0.8 0.56 0.88 0.78 669
WildGraph (1.0) 9.63 0.93 71.08 0.92 2.93 0.81 0.57 0.89 0.80 628
1.5 9.29 0.9 66.5 0.87 2.98 0.8 0.55 0.9 0.81 635
2.0 9.52 0.93 70.15 0.91 3.15 0.81 0.54 0.89 0.8 632
Stork 0.25 7.14 0.77 37.6 0.75 3.44 0.65 0.61 0.82 0.78 407
0.5 6.97 0.8 35.88 0.75 3.29 0.7 0.6 0.83 0.74 447
WildGraph (1.0) 6.61 0.84 35.98 0.81 3.6 0.74 0.59 0.83 0.78 438
1.5 7.07 0.82 37.84 0.77 3.61 0.71 0.58 0.86 0.74 422
2.0 6.84 0.86 36.65 0.78 3.51 0.73 0.59 0.84 0.75 352
@@ 6. Conclusions
In this paper, we study the problem of trajectory generation over a long horizon, with a focus on applications in wildlife in a small data setting. Existing methods fail to consider the underlying network that describes this movement. In order to solve this problem, we introduced a hierarchical prototype network that captures the global movement characteristics and recursively localizes regions. We proposed a novel VRN to capture the transitions and feed them into a latent dictionary. The effective use of the H3 library, and node2vec embeddings has contributed to the success of the WildGraph framework, as measured in the experimental results and ablation study.
This study adds to the body of literature where graph-based solutions can effectively address unique problems. Ours is the first to show how one can generate trajectories for wildlife movement. In our case, the small data nature of wildlife datasets combined with wide-horizon predictions presented significant challenges. Our results demonstrate the effectiveness of our proposed approaches that utilize specially designed encoder-decoder network and work remarkably well.
Furthermore, similar to other generative models, such as GPT, our method treats regions as tokens and trajectories as an ordered list of tokens. With increased training data, it may be possible to fine-tune for downstream applications, such as identifying bird characteristics, like age or type, based on their movement patterns.
Finally, it may be interesting to explore how this approach could be adapted for similar problems such as the trajectory prediction task. Additionally, the simplicity of the approach might generalize well to congruous problems in the vehicle movement or human movement domains.
Appendix A Theoretical Analysis on Probability of Replication
Here, we investigate the probability of WildGraph generating a sequence of regions that replicates a real trajectory in the training set. First, to summarize WildGraph, at each transition step the VRN performs the following:
  1. (1)
    Find the $y$ most probable states.
  2. (2)
    Find the corresponding $y$ node2vec representation vectors.
  3. (3)
    Compute the weighted average for these vectors, i.e. belief vector.
  4. (4)
    Utilize this belief vector as an input to the decoder during the generation.
A simple theoretical analysis of the methods demonstrates an extremely low probability of replication by our approach.
Proposition A.1.
Given h3 regions $h3_{i},h3_{i+1}$, the transition probability $P(h3_{i+1}|h3_{i})$ generated by VRN is $\approx\frac{1}{y}$, where $y$ is the number of top most probable nodes used in the belief vector.
Proof.
Since the belief vector represents the $y$-most probable nodes, it is expected that the next generated node can be one of them. The expected probability of a transition is, therefore, congruous to $\frac{1}{y}$, given $h3_{i+1}=\sum_{i=1}^{y}Prob(h3_{i})\times h3_{emb_{i}}$. ∎
Given training trajectories $T_{train}$ (each of length $m$) that consist of regions $(h3_{1},h3_{2},\ldots)$, the probability of generating a $S$ with an identical ordering of regions is upper-bounded by $(\max Prob(h3_{i},h3_{i+1}))^{m}$, where $Prob(h3_{i},h3_{i+1})$ is the most common transition probability in $T_{train}$. Given $Prob(h3_{i},h3_{i+1})\approx\frac{1}{y}$, the probability of duplicate trajectories is $(\frac{1}{y})^{m}\to 0$ when $y>1$ and $m>>1$.